21. Copper -64 is radioactive and has a half life of 12.7 hours. The sample that initially contains 3mh of Cu-64, How much time will it take for only .5mg to remain. 22. Find the pH of a solution .15M of propionic acid. Ka for the acid is 1.3x10n-5Please show every step!
Answer #1

21.

All radio active decay follows, first order kinetics,

relation between first oreder rate constant and half life time is,

k = 0.693 / t1/2 = 0.693 / 12.7 = 0.0546 hour-1

First oreder integrated rate constant equation is,

k = (1/t)*Ln[A]0/[A]t

t = (1/0.0546) * ln(3.0 / 0.5)

t = 32.8 hours

(22)

According to Ostwalds' theory of weak acids,

[H3O+] = sqrt.(Ka*C) = sqrt.(1.3*10-5*0.15) = 1.40 * 10-3 M

therefore,

pH = - Log[H3O+] = - Log(1.40*10-3) = 2.85